OpenStudy (sloppycanada):

Solve x + y = 5 using matrices. 2x – y = 1 I have the answer, x = 2 and y =3, how do I phrase this as a matrices? (2,3)?

rishavraj (rishavraj):

u know Cramer's Rule??

OpenStudy (sloppycanada):

I don't think that was covered.

rishavraj (rishavraj):

Cramer's Rule. If the equations are \[a_1 x + b_1 y = c_1\] \[a_2 x + b_2 y = c_2\] then the solutions are are \[x = \frac{ D_1 }{ D } ~~~~ and ~~~~~ y = \frac{ D_2 }{ D }\]

rishavraj (rishavraj):

where 

OpenStudy (sloppycanada):

Okay... so I'm phrasing mine wrong?

rishavraj (rishavraj):

hmmm wht do u mean?? u tried it using matrices??

OpenStudy (sloppycanada):

I think so, I'm thinking it's [2 3]

rishavraj (rishavraj):

u hav been askd to solve using matrices ...thts why u need to solve it tht way

OpenStudy (sloppycanada):

but I don't understand how I'm supposed to do that. Just plug in the equations to what you gave me?

rishavraj (rishavraj):

yup....bt use with appropriate signs.....the answers will remain the same....kinda obvious :P

OpenStudy (sloppycanada):